OpenStudy (anonymous):

Solve 2x2 - 8x = -7 using the quadratic formula

OpenStudy (anonymous):

2^2

OpenStudy (callisto):

2x^2 - 8x = -7 2x^2 - 8x +7 = -7 +7 2x2 - 8x +7 = 0 Compare that with ax^2 + bx = c = 0 What is a, b and c?

OpenStudy (anonymous):

2 -8 and 7

OpenStudy (callisto):

Yes, apply the quadratic formula to solve x. Quadratic formula: \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (anonymous):

i did and i got all the way to \[8\pm √64 - 224 all divided by 4\]

OpenStudy (anonymous):

and then i got stuck

OpenStudy (callisto):

What is 4ac ??

OpenStudy (callisto):

where a = 2 and c = 7

OpenStudy (anonymous):

4(2)(7) 8 * 28 right? which then equals 224

OpenStudy (callisto):

No!!! 4ac = 4 (2) (7) = 8 (7) = ...?

OpenStudy (anonymous):

ohhh ok thanks!

OpenStudy (callisto):

Welcome ._.